1. Find the Laplace transform of the function f(t) = 1 + 2t + 3e-3t - 5 sin(4t). Solution:

2. Find the inverse Laplace transform of F(s) = 7+ (8 + 4)(18 - 3s) (s - 3)(s – 1)(s + 4) Solution:
Answer #1

~34 0 Given set=1t2ttzéss Sin (4+) - Taking Lot; . Lifet y =.L{ 14+2lsts +36h ésty -5 ch. Sin (40) = 1 + 2 + 343 - 2016) [: 2At B t = -3 3A+B-4D=6 -4A-12 B + 3D=79 Solving we get A=5, B=-41/1,0=15 Putting there in o and - taking inverne Lit; [s Flod